Solve the following system of algebraic equations.
a. ?=?2 and ?=3?
b. 2?=?2 and ?=4−?2
c. 12?−?=0 and 3?2−4?−8=0
Answer #1

Given y=x², Y=3x 2 Y=x2 Y = 32 o= x² 3x X =3, X=0. put the solutions x=3,a=0 into Y=x² for y=x² substitute x with 3 y = 9. foTherefore, final solutions are (x=3, y=9 x=0, y=0 2) Given 24= 12 Y=4-x² 2y=x? +Y 4-X? 2y+y = x +(4-23 3y = 4 y = 4 3For 12x-1=0, substitate ? Put the solutions y= 4 into ey=x² for 2y=x², substitute y with 4 &(4) = x² x=-256 256 . Therefore,Therefore, the solutions are x=2 (12+8556) 2,9=96+4056 x = 2(12-51 Y=96-406 3